A bottler of drinking water fills plastic bottles with a mean volume of 999 milliliters (mL) and standard deviation 5 mL. The fill volumes are normally distributed. What proportion of bottles have volumes between 992 mL and 998 mL?

1
Expert's answer
2022-01-22T14:55:46-0500

μ=999mL

σ=5

P(992<X<998)="P(\frac{992-999}{5}<Z<\frac{998-999}{5})\\
P(-1.4<Z<-0.2)\\
P(Z<-0.2)-p(Z<-1.4)\\
0.4207-0.0808=0.3399"